Solve each equation
2(5x+58)=10x+4(21+3.5-11)

Answer by checkley75(3666)

2(5x+58)=10x+4(21+3.5-11)
10x+116=10x+4(13.5)
10x-10x=-116+54
0x=-62
no solution for x.